OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y

OpenStudy (anonymous):

are you familiar with the substitution method?

OpenStudy (anonymous):

yeah a little but its hard to understand

OpenStudy (anonymous):

okay well since x = (1-4y) plug (1-4y) into the other equation where the x is. in order to solve for a variable the entire equation has to be in terms of one variable.. so: 2(1-4y)+7y=3 2-8y+7y=3 2-y=3 y=-1 so now knowing y=-1.. assuming i haven't made any mistakes, you can put that into either equation and then solve for x.. hope this helps?! let me know

OpenStudy (anonymous):

wait let me check this.

OpenStudy (anonymous):

if you want to check your answer, you should find that x = 5, y = -1